Focus on the difference between feasible alternatives (Principle 2)l Insulated concrete forms (ICF) can be used as a than in a similar wood-framed building in upstate New York. An ICF home will be approximately 10% more expensi York and costing $210 per month to heat and cool, how many months does it take for a 2,100 ft ICF home to pay It will take months to pay back the extra ICF construction cost through monthly energy savings. (Round to the
n be used as a substitute for conventional wood framing in building construction. Heating and cooling bills will be about 60% less more expensive to construct than a wood-framed home. For a typical 2,100 ft home costing $150 ft to construct in upstate New nome to pay back its extra construction cost? ound to the nearest whole number.)
Answer #1

Here, we are comparing two alternatives. The alternatives are 1> Wood Frame 2> ICF

Bill is lesser for the ICF but the construction cost of ICF is high.

The ICF costs $150x2100=$315,000 for a typical home which is 10% higher. So, it costs $31,500 more. This amount is to be recovered.

The monthly bill is $210 which is 60% lower. Thus, x=210/0.6=$350 will be the monthly bill for the wood frame. It is $140 more.

So, it will take 31500/140=225 months to recover the cost.